OpenStudy (anonymous):

It is a question about space vector. At the time t=0, a particle start at point (1,2,3), it along a straight line to point (4,1,4). The speed is 2 at the point (1,2,3) and it has a constant acceleration which is (3i-j+k). Find the location of particle r(t) at time t.

OpenStudy (anonymous):

s = ut + 0.5 a t2?

OpenStudy (anonymous):

u got everythin...............so substitute

OpenStudy (anonymous):

s = s2-s1

OpenStudy (anonymous):

well i translate it from a chinese book...maybe it's my reason to let you hard understand it the answer it gave is far than yours :(

OpenStudy (anonymous):

oh.... :P but this is the way :P u got speed.....................u got to put velocity vector.... u kno a unit vector.....parallel to that

OpenStudy (anonymous):

i was thinking of that approach too^ and think the answer would be (2/11)(3,-1,1)t + 0.5t(-3,-1,1)

OpenStudy (anonymous):

um i think u shud add (1,2,3)

OpenStudy (anonymous):

oh yah u're right (1,2,3)+(2/11)(3,-1,1)t + 0.5t(-3,-1,1)

OpenStudy (anonymous):

@AndrewNJ dude is that it?

OpenStudy (anonymous):

the answer is (0.5t^2+2t/(Square root of 11))(3i-j+k)+(i+2j+3k) terrible internet, i spent 20 min refreshing this page!

OpenStudy (anonymous):

@Tushara @A.Avinash_Goutham upper is the answer

OpenStudy (anonymous):

my bad it shud hav been sqrt(11)

OpenStudy (anonymous):

ok so dat is the same as (i+2j+3k)+(0.5t^2)(-3i-j+k)+(2/sqrt11)(-3i-j+k) do you understand how to get to dat answer?

OpenStudy (anonymous):

i only feel puzzled about where to get 2t/sqrt11

OpenStudy (anonymous):

other part is ok

OpenStudy (anonymous):

well u're finding the unit vector in da direction of the velocity: (4,1,4)-(1,2,3)=(3,-1,1) sqrt(3^2+(-1)^2+1^2) = sqrt(11) hence unit vector= (1/sqrt11)(3,-1,1) which is direction vector and you have to multiply it by the speed which is 2 therefore you get (2/sqrt11)(3,-1,1)

OpenStudy (anonymous):

and since you have to multiply velocity by time, you get (2t/sqrt11)(3,-1,1)

OpenStudy (anonymous):

ya... got it thanks!!